# MapFormer: Self-Supervised Learning of Cognitive Maps with Input-Dependent Positional Embeddings

### v2

@@ 7 Limitations
We highlight several limitations of our work that we left for future explorations. First, in order to compute path integration, we sum the rotation angles along the temporal dimension, meaning that our formalism only applies to causal transformers. Further work is needed to generalize our method to non-causal problems such as visual processing, but the same methods used to adapt SSMs to vision zhu2024visionmambaefficientvisual could be used. Second, we did not compare MapFormers to PoPE gopalakrishnan2025decouplingwhatwherepolar, but since it relies on a mechanism similar to CoPE, we expect it to behave similarly. Third, we did not try to scale our method to larger model sizes and larger datasets, but since MapFormers require little extra compute compared to baseline transformers, computation time should not be a bottleneck. Fourth, our results only confirm the superiority of EM models on recall tasks compared to their WM counterparts, but further experiments should compare these two models on more complex tasks. Finally, in our models, for numerical efficiency, we only use block diagonal matrices, without orthogonal matrix $M$ used in eq. 2. However, real neurons cannot have a negative firing rate, and because of that, position $p_{t}$ must always be positive, which is impossible with a block diagonal rotation matrix. Matrix $M$ mixes frequencies together to make the code non negative dorrell2023actionableneuralrepresentationsgrid, which creates a fundamental difference between our models and biological systems. Non-negativity and bounded activity can also result in disentangled representations whittington2023disentanglementbiologicalconstraintstheory, which could help build even more robust and interpretable cognitive maps in the future.
8 Conclusion
We introduce simple theoretical arguments from Lie-group theory to analyze transformers and SSMs, allowing us to characterize the type of structures that these models are able to learn, and show that diagonal matrices used in current SSMs are not expressive enough to learn spatial cognitive maps. From this, we introduce MapFormer, which generalizes the Transformer architecture with input-dependent transformations, and unify relative and absolute positional embeddings with computational models of working and episodic memory, respectively. Our models leverage the parallel sequential processing ability of Transformers to compute path integration in parallel, mandatory for larger-scale applications, and with a minimal structural bias, learn a cognitive map within their positional embeddings via prediction of upcoming observations. MapFormers are, to the best of our knowledge, the first to learn such maps without any supervision and learn to represent actions as invertible matrices, facilitating generalization by ensuring a consistent application of these actions across samples, while increasing explainability via a direct access to the sequence of actions taken by the model.
This work has implications in both neuroscience and AI: for the former, it explains the neural mechanisms giving rise to cognitive maps; in machine learning, this work allows to learn structure and relations at scale, creating smaller, more robust and interpretable models.
@@ 9 Acknowledgments
10 Appendix
@@
10.1 MAmPa: Mamba with skew-symmetric block-diagonal matrices
As explained in sec.3.4, one only needs to modify the recurrent matrix $A:=S=-S^{\top}$ of selective SSMs to be block-diagonal skew-symmetric in order to learn spatial cognitive maps. However, the reason why people chose diagonal matrices in Mamba is to dramatically reduce the inference speed of these models, and make them as close as possible to parallelized. In Mamba, an efficient implementation of the parallel scan is used, leveraging the fact that $A$ is diagonal and can therefore be treated as a vector. In Mamba2, the authors go even further by choosing $A:=a\in\mathbb{R}$ to be a scalar, which allows them to broadcast it easily in the attention matrix and increase the model’s dimension (eq. 16), showing that recent attempts at scaling SSMs mainly rely on decreasing the selective weights expressivity, and therefore learning of cognitive maps at scale.
@@ 10.1 MAmPa: Mamba with skew-symmetric block-diagonal matrices
However, as we have demonstrated in our experiments, diagonal matrices are not expressive enough to learn spatial cognitive maps, nor any other higher dimensional map. We implemented a parallel scan on square matrices to test if MAmPa, an extended version of Mamba with $2\times 2$ block-diagonal matrices can learn spatial maps like our theory predicts. Because the parallel scan on square matrices dramatically increases the computational load, we were forced to keep the inner state small ($N=16$) and only trained models on sequences of size $l=16$ since we did not take the time to implement a custom parallel scan using the efficient implementation of matrix-vector product for rotations used in RoPE (eq. 18), but we expect it to be highly beneficial to lean cognitive maps at scale with Mamba models. We report the results in tab. 3, where OOD dense and OOD sparse use sequences of size $8$ and $32$ respectively.
@@
10.2 Frequency initialization

### v3

@@ 7 Limitations
We highlight several limitations of our work that we leave for future explorations. First, in order to compute path integration, we sum the rotation angles along the temporal dimension, meaning that our formalism only applies to causal transformers. Further work is needed to generalize our method to non-causal problems such as visual processing, but the same methods used to adapt SSMs to vision (Zhu et al., 2024) could be used. Second, we did not try to scale our method to larger model sizes and larger datasets, but since MapFormers require little extra compute compared to baseline transformers, processing speed should not be a bottleneck. Finally, our results only confirm the superiority of EM models on recall tasks compared to their WM counterparts, but further experiments should compare these two models on more complex tasks such as reasoning.
Impact Statement
The scalability and structural flexibility of MapFormers suggest they could serve as generic architectures for a variety of complex tasks beyond spatial navigation, including language modeling, vision, and protein folding (siméoni2025dinov3; Jumper et al., 2021). By framing the learning of cognitive maps through a Lie-group theoretic lens, our work provides a path toward disentangling abstract structures from raw content. Furthermore, while this work focused on rotation matrices for efficiency, the ability of MapFormers to learn more complex group structures opens new doors for AI. It allows for the exploration of structures where irreducible representations are unknown, such as rotations in high-dimensional spaces or non-commutative groups. This offers a theoretical framework for understanding how biological and artificial agents might represent complex relational data, potentially leading to more robust, structure-aware AI systems that mirror the efficiency of human cognition.
@@ 8 Acknowledgments
Appendix A Appendix
A.1 Action-Dependent Matrices Form a Lie Group
The key idea to learn a cognitive map is to move beyond fixed matrices and learn action-dependent matrices, specific to each input action. In (Dorrell et al., 2023), the authors introduce actionable representations using group and representation theory, as a formalism to map an input action $a\in\mathbb{R}^{r}$ (e.g. in 2D, $r=2$ and ‘move right’) to a neural transformation $W_{a}\in M_{n}(\mathbb{R})$ (e.g., the corresponding weight matrix of a neural network that implements the effect of this action on the represented position in space $p$):
$\phi:\mathbb{R}^{r}$ $\to M_{n}(\mathbb{R})$ (3)
$a$ $\mapsto W_{a}$
The insight that action-dependent matrices need to form a group, and in particular a Lie group over $\mathbb{R}$, allows the use of results from group and representation theory to understand what are the possible constraints on the form of $W_{a}$’s (Dorrell et al., 2023). Specifically, the Peter–Weyl theorem (Knapp and Knapp, 1996) states that any compact group $\mathbf{G}$ can be represented in $M_{n}(\mathbb{R})$ as:
where $\omega_{k},\Delta_{t_{k}}\in\mathbb{R}^{r},r=1,2$ are the angular frequencies and rotation duration, respectively, the former representing movement at different scales. That is, in this example, each $I_{k}$ performs a rotation of a subspace of the neural space. The matrix $M$ then performs a change of basis on this result, assuming $M$ is not the identity.
A.2 Path Integration Can Be More Easily Computed in the Lie-Algebra Space
To learn a cognitive map, we saw that it is desirable to require action-dependent matrices to form a Lie group $\mathbf{G}$. A Lie group is a group $G$ that is also a smooth manifold, such that the group operations (multiplication and inversion) are smooth maps. For example, the group of 2D rotations $\mathrm{SO}(2)$ is a Lie group. To build a neural model that can learn cognitive maps, the goal would be to learn action-dependent matrices that form a Lie group.
To further characterize the group of action-dependent matrices, it is often simpler to analyze the corresponding Lie algebra. The associated Lie algebra $\mathfrak{g}$ of a Lie group $\mathbf{G}$ is the tangent vector space of $\mathbf{G}$ at the identity element. Lie-group theory shows the existence of an exponential map between the Lie algebra and the group $\exp(\cdot):\mathfrak{g}\to\mathbf{G}$, and in the case of a compact group, where the exponential map is surjective, any element $W_{a}\in\mathbf{G}$ of the group can be expressed as the exponential of an algebra element $A$, $W_{a}=\exp(A)\in\mathbf{G}$. This is useful since studying the structure of Lie groups (which involves, e.g., matrix multiplication) is often simpler from the point of view of the corresponding Lie algebra (which involves addition operations).
For the case of cognitive maps, we will think of elements of the Lie algebra $A$’s as action representations, and the corresponding group elements (expressed via the exponential map) as the action-dependent matrices $W_{a}$’s, which implement these actions in synaptic weights. The exponential map provides the link between the vector space of actions (the Lie algebra) and the space of weight matrices (the Lie group). Specifically, given the property of the exponential map (if the $A_{i}$’s commute, see sec. B.2 for further details): $\exp\left(\sum_{i}A_{i}\right)=\prod_{i}\exp(A_{i})$, the exponential mapping translates the additive structure of the Lie algebra (cumulation of actions in this vector space - e.g, ‘move right’ then ‘move up’) into the multiplicative structure of the Lie group (multiplication of the corresponding action-dependent weight matrices). This is useful for path integration, since the cumulative effect of taking actions $A_{1},...,A_{n}$ can be translated to the final action-dependent weight matrix $W_{a}=\prod_{i=1}^{n}\exp(A_{i})$. As we will see, the exponential map allows in some cases to compute path integration in parallel, in the vector space of the Lie algebra, by first computing the cumulative summation of subsequent actions and only then computing the exponent of the result.
For example, for spatial problems, such as navigation on a circle or a torus—and more generally, any 1-dimensional compact group, the Lie group of interest is $\mathrm{SO}(2)$ - the group of all 2D-rotations.
The corresponding Lie algebra of $\mathrm{SO}(2)$, denoted $\mathfrak{so}(2)$, consists of all $2\times 2$ real matrices $A$ that are skew-symmetric ($A^{\top}=-A$). The Lie algebra $\mathfrak{so}(2)$ is a one-dimensional vector space, spanned by a single basis skew-symmetric matrix $S:=-S^{\top}$ (2D-rotations have a single degree of freedom). Thus, all elements of the Lie algebra can be expressed by a single scalar (the angular velocity) $\omega$ as $A=\omega S$. A finite rotation $R_{\theta}$ can then be seen as generated by integrating the infinitesimal angular velocities $\omega$ over a duration $\Delta_{t}$:
Intuitively, the skew-symmetric matrix $A=\omega S$ defines the instantaneous action direction $S$ and constant angular velocity $\omega$. The finite rotation $R_{\theta}\in\mathrm{SO}(2)$ is obtained by the exponential map, which performs a continuous integration of this velocity. Thus, the expression $R_{\theta}=\exp(\Delta_{t}A)$ shows that the final rotation yields a total angle of $\theta=\omega\Delta_{t}$, which defines the smooth path (geodesic) on the group manifold.
More generally, given the group of action-dependent matrices $W_{a}$’s (which, as we saw, needs to be a subgroup of the general linear group $\mathrm{GL}(n,\mathbb{R})$), elements of the algebra, $A(t)$, can be expressed by a linear combination of generators $A_{i}$’s scaled by parameters $t_{i}$’s (the ‘integration times’, which can be seen as coordinates along the different directions defined by the generators). If surjective, all group elements $W_{a}$’s can then be expressed using the exponential map (fig.2):
For the special case where the generators $A_{i}$ commute (i.e., the Lie bracket is zero), as in the simple $\mathrm{SO}(2)$ example, the exponential of the sum equals the product of the exponentials: $W_{a}=\prod_{i=1}^{K}\exp(t_{i}A_{i})$. This shows again the benefit of the Lie algebra: it transforms the non-linear composition (multiplication) of the group into linear path integration (addition) within the algebra, thereby facilitating efficient learning and parallelization in the model.
@@ A.2 Path Integration Can Be More Easily Computed in the Lie-Algebra Space
A.3 MAmPa: Learning Cognitive Maps with block-diagonal Mamba Models
In this section, we propose a Lie-theoretic interpretation of State-Space Models (SSMs), which provides an interpretation of modern SSMs as models that implicitly perform path integration and can thus learn cognitive maps under certain conditions. This framework reveals a substantial limitation of current state-of-the-art SSMs, such as Mamba and Mamba2 (Gu and Dao, 2024; Dao and Gu, 2024), to meet these conditions for learning cognitive maps. We thus propose a simple modification of the Mamba architecture, which we call MAmPa, for learning cognitive Maps with Mamba models. The link established in (Dao and Gu, 2024) between SSMs and Transformers is what will later allow us to understand the links between absolute and relative positional encoding with models of episodic and working memory, respectively.
Recently, structured state-space models (S4) (Gu et al., 2022) have gained popularity over Transformers as they allow parallelized training of sequences while keeping a linear inference cost. These models are inspired by continuous state space models, where a continuous signal $x(t)\in\mathbb{R}$ is mapped to $y(t)\in\mathbb{R}$ via a hidden variable $h\in\mathbb{R}^{n}$:
$h^{\prime}(t)$ $=Ah(t)+Bx(t)$ (9)
$y(t)$ $=Ch(t)$
$h_{t+1}$ $=\overline{A}h_{t}+\overline{B}x_{t+1}$ (10)
$y_{t+1}$ $=Ch_{t+1}$
where $\overline{A}=\exp(\Delta\cdot A)$ and $\overline{B}=(\Delta\cdot A)^{-1}\left(\exp(\Delta\cdot A)-I\right)\Delta B$. This discretization step is fundamental, as it explains how the exponential map can arise from the integration of infinitesimal recurrent transformations over a time interval $\Delta\in\mathbb{R}$. The fact that recurrent computations are define with a single synaptic weight matrix $A$ implies that this exponential map can only generate 1-dimensional Lie-algebras, hence commutative groups.
$\overline{A}_{t}:=\exp(\Delta_{t}A)$ (11)
The exponential relationship between $A$ (the continuous domain matrix) and $\overline{A}$ (the discrete domain matrix) directly parallels the relationship between the element $A$ of the Lie algebra and the weight matrix $W_{a}$ in the corresponding Lie group. In the rotation example, since the Lie-algebra $\mathfrak{so}(2)$ of skew-symmetric matrices gives rise to the Lie-group of rotation matrices $\mathrm{SO}(2)$, $A:=\omega S$ is the infinitesimal generator and $W_{a}$ is the finite rotation over duration $\Delta_{t}$: $W_{a}=\exp(\Delta_{t}A)$. By identifying the discrete-time SSM matrix $\overline{A}$ with the action-dependent matrix $W_{a}$, we propose a new theoretical framework to analyze the capacity of Mamba to learn action-dependent matrices and cognitive maps. In this framework, Mamba can be interpreted as learning action-dependent matrices $\overline{A}_{t}$, which are dynamically created via an exponential map. As we saw, the exponential map integrates infinitesimal steps via $A$ (the direction and velocity in the Lie algebra space) over a varying time interval $\Delta_{t}$, giving a finite transformation $\overline{A}$ (the element of the Lie group). In other words, Mamba implicitly computes path integration of action $A$ over a duration defined by the learned parameter $\Delta_{t}$.
Given this theoretical framework, it is evident that current versions of Mamba are incapable of learning cognitive maps that require, for example, learning rotations. In Mamba, $A$ was chosen as diagonal, to reduce computational cost, but a diagonal action-dependent matrix cannot learn rotations. This is since $\overline{A}$ needs to be similar to a block-diagonal matrix, with blocks of size 2 (eq. 4 & 5). Mamba was shown to be able to learn gating problems, which increases the performance of Mamba on various structure-dependent tasks by ignoring irrelevant information, however, Mamba with a diagonal constraint is not expressive enough to represent spatial cognitive maps as Lie-group theory shows. In this work, we modify $A$ in Mamba to be block-diagonal and, in particular, to blocks of size 2 (for learning rotations) and show that this modification increases the capacity of the model to learn cognitive maps.
A.4 MAmPa as a Model of Episodic Memory (EM) and Working Memory (WM)
Figure 5: An illustration of the conceptual framework of action-dependent matrices, and the differences between Episodic Memory (EM) and Working Memory (WM) models (Whittington et al., 2025). (a) Episodic Memory (EM) model: Positions (or ‘states’, more generally) in a cognitive map $p_{t}$’s (in blue), at times $t$’s, are encoded in a separate neural space by a dedicated neural population, which does not directly interact with the input $x_{t}$ (red), being updated via actions only. Actions $a_{t}$’s, when they occur in the input, can update the position via input-dependent matrices $W_{a_{t}}$ (green): $p_{t}=W_{a_{t}}p_{t-1}$. $W_{a}$’s are input-dependent matrices and are computed through an exponential map, which maps between abstract representations of actions $A_{t}:=\Delta_{t}\omega S=\theta_{t}S$ (in the Lie-Algebra) and the corresponding weight matrix $W_{a_{t}}:=\exp(A_{t})=R_{\theta_{t}}$ (in the corresponding Lie group; section A.2). The exponential mapping is computed based on the integration time $\Delta_{t}$ of the action (light green) – roughly, how long the synaptic weights $S$ need to be applied. Finally, the conjunction (purple) of the content (what) and the structure (where) can be computed. (b) Working Memory (WM) model: In contrast to the EM model, which factorizes content (red) and position (blue) in two separate neural spaces, MapFormers-WM represent position implicitly. Actions in the WM model directly update the conjunction of what and where (purple). The conjunction states $h_{t}$’s in the WM model are thus smaller by a quadratic factor compared to the EM model.
In (Whittington et al., 2025), the authors introduce two related models of episodic (EM) and working memory (WM) that both learn cognitive maps using RNNs with input-dependent matrices. Here, we explain why selective, discrete State-Space Models (SSMs) can be viewed as EM and WM models. First, note that discrete SSMs can be regarded as RNNs with no activation function. Second, selective SSMs learn input-dependent matrices via an exponential map (eq. 10), meaning that their inner state captures the task’s latent structure, as in EM/WM models. Therefore, EM and WM models can naturally be regarded as selective SSMs.
SSM as a model of episodic memory
The EM-RNN model behaves exactly like TEM (Whittington et al., 2019), factorizing structure and content in two distinct neural populations, but only updates its position with an input-dependent RNN $p_{t}:=W_{a_{t}}p_{t-1}$. It then queries a Hopfield memory module (Ramsauer et al., 2021) with its updated position to retrieve passed observations. The RNN in TEM can be seen as an SSM with subtle differences: (1) it uses no external input, i.e. $\forall t,x_{t}=0$; (2) a learnable embedding $p_{*}$ representing the cognitive map’s zero coordinate must be introduced to avoid the state being always null; (3) SSM’s inner state must be projected back to its final state ($y_{t}=Ch_{t}$, eq.10), which implies in this case that the final position $q_{t}:=Cp_{t}$ can be obtained with an optional learnable projection $C$; (4) as in TEM, the memory module is queried with the conjunction $g_{t}$ of what has been seen and where it was observed (fig. 2(a)):
$q_{t}$ $=Cp_{t}=CR_{\theta_{0\to t}}p_{*}$
$g_{t}$ $=\text{vec}(x_{t}^{\top}.q_{t})$
Setting $C:=\mathbf{1}_{n}$ implies that $q_{t}=p_{t}$, which is analogous to TEM/EM RNNs. Moreover, compared to RNN based models, this formulation allows to compute path integration directly in the Lie algebra ($\omega S\sum_{0}^{t}\Delta_{s}$) and avoid sequential matrix multiplication.
SSM as a model of working memory
WM-RNNs directly updates the internal state with input-dependent matrices, which is precisely what selective SSMs do. As such, current selective SSMs with the appropriate matrix structure (skew-symmetric in the case of navigation, eq. 5) naturally model working memory (fig. 2(b)).
In the previous section, analyzing SSMs through Lie-group theory has allowed us to understand their current limitations, and that one only needs to switch to block-diagonal matrices to theoretically be able to learn a cognitive map. However, previous work decided to keep diagonal matrices because it dramatically reduces the computational load. By treating $A$ as a vector, one can avoid computing a matrix exponential and also dramatically reduce the memory footprint of a parallel scan, by computing vector-vector multiplications instead of matrix-vector’s. In Mamba2 (Dao and Gu, 2024), it was further reduced to a scalar $A:=a\in\mathbb{R}$, which allows them to implement a fully parallelized version of selective SSMs but looses the ability to learn cognitive maps, as we have seen above. This highlights a fatal limitations of parallel-selective SSMs: they can be parallelized at the cost of loosing any hope of learning a cognitive map (see sec. A.5 for further details). This is the reason why we decided to move away from SSMs and adapt Transformers to learn cognitive maps, as models of episodic (MapFormer-EM) and working memory (MapFormer-WM).
A.5 MAmPa: Mamba with skew-symmetric block-diagonal matrices
@@
As explained in sec.A.3, one only needs to modify the recurrent matrix $A:=S=-S^{\top}$ of selective SSMs to be block-diagonal skew-symmetric in order to learn spatial cognitive maps. We implemented a parallel scan on square matrices to test if MAmPa, an extended version of Mamba with $2\times 2$ block-diagonal matrices can learn spatial maps like our theory predicts. Because the parallel scan on square matrices dramatically increases the computational load, we kept the inner state small ($N=16$) and only trained models on sequences of size $l=16$ since we did not take the time to implement a custom parallel scan using the efficient implementation of matrix-vector product for rotations used in RoPE (eq. 16), but we expect it to be highly beneficial to lean cognitive maps at scale with Mamba models. We report the results in tab. 3, where OOD dense and OOD sparse use sequences of size $8$ and $32$ respectively.
@@
A.6 Relationship between MapFormers and MAmPa: Linear MapWM are selective SSMs
In sec. 4.1, we have introduced the exponential map in Transformers (via the rotation matrices) to allow them to build input-dependent matrices and learn cognitive maps. The Transformer formalism cannot naturally explain how an exponential map emerges like in SSMs, but in Mamba2 (Dao and Gu, 2024), the authors show that linear transformers (Wang et al., 2020) are actually SSMs. Starting from equation 10:
$h_{t}$ $=A_{t}h_{t-1}+B_{t}x_{t}$ (13)
after multiplying by $C_{t}$, we get the final equation:
$y_{t}=\sum_{s=0}^{t}C_{t}^{T}A^{\times}_{s\to t}B_{s}x_{s}$ (14)
$\Rightarrow$ $y=Mx;\quad m_{ij}=C_{j}^{T}A_{j}.A_{i+1}B_{i}$
If we choose matrix $A:=S$ to be a $2\times 2$ block-diagonal skew-symmetric matrix and $R_{t}:=\exp(\theta_{t}S)=\exp(\Delta_{t}\omega S)$ as in equation 7, we retrieve the definition of RoPE (Su et al., 2023):
$m_{ij}$ $=q_{j}R^{\times}_{i\to j}k_{i}$ (15)
$=q^{T}_{j}R^{\times}_{0\to j}\left(R^{\times}_{0\to i}\right)^{-1}k_{i}$
$=\text{RoPE}(q_{j},k_{j})$
This proves that linear RoPE transformers are SSMs with a $\mathbf{2\times 2}$ skew-symmetric recurrent matrix, where $\theta_{t}$ controls the angle of rotation. In Mamba2, the authors define $A:=a\in\mathbb{R}$ as a scalar, giving $M_{ij}=\exp(-a\sum_{t=i+1}^{j}\Delta_{t})q_{j}^{\top}k_{i}$, which is analogous to the cumulative gating of CoPE (Contextualized Positional Encoding) (Golovneva et al., 2024), with a multiplicative and not additive binding operation. CoPE is an enhanced Transformer architecture with context-dependent positional encoding, that like Mamba compresses the relative key-query distances with a cumulative gating mechanism, and in both models, input selectivity allows them to focus or ignore certain tokens and solve structure-dependent tasks, but should not allow them to learn spatial cognitive maps.
This extends the unification of transformers with models of episodic memory (EM) started in TEM-t (Whittington et al., 2022), showing that the difference between absolute and relative positional encoding can be better understood via the SSM formulation. EM models use action-dependent matrices on a dedicated neural population (absolute positional encoding), before querying an external memory module, i.e. modeling memories in synaptic activity, while WM models directly update the whole neural activity with action-dependent matrices, directly
A.7 Implementation details
Rotations can be ensured by enforcing $S:=\overline{S}-\overline{S}^{\top}$ to be skew-symmetric, but since a single generator matrix $S$ can only generate rotations along a single axis (see sec.B.2.2 for further details), choosing blocks of size $b>2$ is useless as it increases the computational load while effectively reducing the amount of rotations being performed. Using blocks of size $2$ also allows us to use the explicit irreps representation of $2\times 2$ rotations and avoid computing a matrix exponential (eq.5). This also allows, as in RoPE (Su et al., 2023), to rewrite matrix multiplication of $R_{\theta}x$ in MapFormers as:
leaving only angular velocities $\omega\in\mathbb{R}^{n_{b}}$ (see sec. A.8 for a discussion about their initialization) as learnable parameters instead of matrix $S\in\mathbb{R}^{n_{b}\times b\times b}$.
The actual rotation angle $\theta_{t}=\omega\Delta_{t}$ is dynamically computed from the input features $X$ using a low-rank projection $W_{\Delta}=W_{\Delta}^{out}W_{\Delta}^{in}$. The goal of this two-step process is to ensure the rotation relies only on the most salient action features of the input. First, the internal projection $W_{\Delta}^{in}\in\mathbb{R}^{d\times r}$ ($r\ll d$) maps the high-dimensional input $X$ to a low-dimensional $\Delta_{t}^{in}\in\mathbb{R}^{t\times r}$ (for instance, in a 2D environment, $\Delta_{t}^{in}$ could be the 2D movement vector $\Delta_{t_{k}}\in\mathbb{R}^{r}$ in eq. 5, where $r=2$). Second, this extracted action is projected by $W_{\Delta}^{out}$ into the final set of path increments $\Delta_{t}\in\mathbb{R}^{t\times n_{h}\times n_{b}}$, which controls the rotation angle for all attention heads ($n_{h}$) and all diagonal blocks ($n_{b}$), where $n_{b}=d_{h}/b$ is the number of blocks per head dimension.
Finally, compared to TEM and TEM-t that use a single neural population $p_{t}$ to encode position, our MapFormers use two separate initial vectors $k_{0}^{p}$ and $q_{0}^{p}$. This distinction is optional and mimics the formalism of EM-SSMs defined in eq. 12, meaning that we could set $k_{0}^{p}=q_{0}^{p}=p_{0}$ without loss of generality. However, we suspect this separation to be beneficial because it would create sparser attention values. Indeed, our positions $p_{t}$ vary smoothly through path-integration, implying that nearby positions would not be strongly separated, but introducing distinct neural populations for keys and queries could allow the model to find the optimal balance between separation/matching of distinct/similar positions.
Recall that the conjunctive query $g^{q}_{i}:=\text{vec}({q^{x}_{j}}^{\top}.q^{p}_{i})$, the conjunction of what has been seen and where. Since $\forall i,j:\langle g^{q}_{i},g^{k}_{j}\rangle=\langle{q^{x}_{j}}^{\top}.q^{p}_{i},{k^{x}_{i}}^{\top}.k^{p}_{j}\rangle=\langle{q^{x}_{j}},k_{j}^{x}\rangle\langle q^{p}_{i},k^{p}_{j}\rangle$, we can avoid the outer products $Q_{X}^{\top}Q_{P}$ and $K_{X}^{\top}K_{P}$ by computing content and position attentions $A_{X}=\textbf{Att}(Q_{X},K_{X})$ and $A_{P}=\textbf{Att}(Q_{P},K_{P})$ separately (still in parallel by concatenating the heads), such that $A_{P}$ acts as an attention mask on $A_{X}$
A.8 Angular velocities initialization